# Supersaturation of Tree Posets

### v1

@@ 3. $(\hat{c})$
Note that the lemmas as stated in [5] hold for $\mathcal{F}$ lying in $\{F:||F|-n/2|\leq 2\sqrt{n\log n}\}$, but since ~$\widetilde{\mathcal{B}}_{n}\subseteq\{F:||F|-n/2|\leq 2\sqrt{n\log n}\}$ their lemma continues to hold for ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ as stated.
Let $\mathcal{M}$ be a family of $q$-marked chains. Let $i\in[q]$, $F\in\mathcal{B}_{n}$. Recall that for each $\chi\in\mathcal{C}$, we let $\mathcal{M}(\chi,F,i)$ denote the set of all $(\chi,Q)\in\mathcal{M}$ such that $F$ is the $i$-th member of $Q$, and we let $\mathcal{M}(F,i)=\bigcup_{\chi}{\mathcal{M}(\chi,F,i)}$. Recall also that for each $s\in[q]$,
$\mathcal{L}^{s}(\mathcal{M})=\{D\in\mathcal{B}_{n}:D\text{ is the }s\text{th member of a }q\text{-chain in }\mathcal{M}\}.$
For $q\geq 2$, we say that a $q$-marked chain family $\mathcal{M}$ is $\ell$-gapped if for every $(\chi,Q)\in\mathcal{M}$ and $F,G\in Q$ distinct with $F\subsetneq G$, we have $|G-F|\geq\ell$.
@@
Let $i,s\in[q]$ such that $i\neq s$ and fix $\ell\geq 1$. For every $\delta>0$, there exists $\gamma=\gamma(\ell,q)$ such that the following holds. Let $\mathcal{M}$ be a $\ell$-gapped family of $q$-marked chains with markers from ~$\widetilde{\mathcal{B}}_{n}$ and let $F\in\mathcal{L}^{i}(\mathcal{M})$. Suppose $F$ is $(i,s,\gamma)$-bad with respect to $\mathcal{M}$ and that $n$ is sufficiently large.
@@
$\operatorname{Prob}[\chi\cap\mathcal{W}_{2}\neq\emptyset|F\in\chi]$ $\leq\sum_{D\in\mathcal{W}_{2}}\frac{1}{\binom{|F|}{|F-D|}}$ (By Equation (6))
@@
$\leq\frac{\delta}{2}$
by choice of $\gamma=\frac{\delta}{2(3\ell q)^{\ell q}}$. On the other hand by Proposition 6.1,
@@
$\operatorname{Prob}[\chi\cap\mathcal{W}\neq\emptyset|F\in\chi]\leq\delta$
@@
Given a poset $P$ and an integer $q$, we say that a map $I:P\to[q]$ is admissible if $x<y$ in $P$ implies $I(x)>I(y)$. With this definition we can state our main lower bound result.
@@
If $P$ is a tree poset of height $k\leq q$ and if $I:P\rightarrow[q]$ is admissible, then for any ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ of size $(q-1+\varepsilon){n\choose n/2}$, we have that the number of induced copies of $P$ in $\mathcal{F}$ is at least
$\Omega\left(\prod_{xy\in H(P)}n^{|I(y)-I(x)|}\cdot{n\choose n/2}\right),$
where here $H(P)$ denotes the Hasse diagram of $P$.
@@
For some slight ease of notation, we prove this result only for tree posets $P$ with at least three members, the case of $P$ having only two members being already addressed by Kleitman’s result for 2-chains [17].
Because $P$ is a tree poset, there exists an ordering $v_{1},v_{2},\dots v_{|P|}$ of the members of $P$ such that every member $v_{j}$ with $j\geq 2$ has exactly one neighbor $v_{j^{\prime}}$ with $j^{\prime}<j$ in the Hasse diagram of $P$, and we call this $v_{j^{\prime}}$ the parent of $v_{j}$.
For this proof, we will need to work with a “nice” subset $\mathcal{F}^{\prime}\subseteq\mathcal{F}$ with large Lubell weight. This set will be defined in one of two ways depending on how many members of $\mathcal{F}$ are close to the middle level of $\mathcal{B}_{n}$.
Case 1: There are at least $(q-1+\frac{\varepsilon}{2})\binom{n}{n/2}$ members $F$ of $\mathcal{F}$ that satisfy:
$\binom{n}{|F|}\geq\frac{\varepsilon}{4q}\binom{n}{n/2}.$
In this case we let
:=$\mathcal{F}^{\prime}:=\left\{F\in\mathcal{F}:\binom{n}{|F|}\geq\frac{\varepsilon}{4q}\binom{n}{n/2}\right\}$
and observe that by 6,
$\mu(\mathcal{F}^{\prime})=\sum_{F\in\mathcal{F}^{\prime}}\frac{1}{\binom{n}{|F|}}\geq\frac{|\mathcal{F}^{\prime}|}{{n\choose n/2}}\geq q-1+\frac{\varepsilon}{2}.$
Case 2: There are least $\frac{\varepsilon}{2}\binom{n}{n/2}$ members $F$ of $\mathcal{F}$ that satisfy
$\binom{n}{|F|}\leq\frac{\varepsilon}{4q}\binom{n}{n/2}.$
Let ′′=$\mathcal{F}^{\prime\prime}=\{F\in\mathcal{F}:\binom{n}{|F|}\leq\frac{\varepsilon}{4q}\binom{n}{n/2}\}$. We have
$\mu(\mathcal{F}^{\prime\prime})=\sum_{F\in\mathcal{F}^{\prime\prime}}\frac{1}{\binom{n}{|F|}}\geq|\mathcal{F}^{\prime\prime}|\frac{4q\varepsilon}{\binom{n}{n/2}}\geq\frac{\varepsilon}{2}\frac{4q}{\varepsilon}\geq 2q.$
Let $\mathcal{F}^{\prime\prime}_{O}=\{F\in\mathcal{F}^{\prime\prime}:|F|\text{ is odd }\}$ and $\mathcal{F}^{\prime\prime}_{E}=\{F\in\mathcal{F}^{\prime\prime}:|F|\text{ is even}\}$. Then either $\mu(\mathcal{F}^{\prime\prime}_{O})\geq q$ or $\mu(\mathcal{F}^{\prime\prime}_{E})\geq q$, and we let $\mathcal{F}^{\prime}$ be the one between $\mathcal{F}^{\prime\prime}_{O}$ and $\mathcal{F}^{\prime\prime}_{E}$ with $\mu(\mathcal{F}^{\prime})\geq q$.
In either of these two cases, we have $\mu(\mathcal{F}^{\prime})\geq q-1+\varepsilon/2$, so we can apply Theorem 5.9 to obtain a nested family of $q$-marked chains $\mathcal{M}^{|P|}\subseteq\cdots\subseteq\mathcal{M}^{0}$ with markers in $\mathcal{F}^{\prime}$ satisfying the two properties of Theorem 5.9. Fix $\gamma$ to be the $\gamma(2,q)$ returned by Lemma 6.3. Moreover, if we let $\ell:=1$ in Case 1 and $\ell:=2$ in Case 2, then we observe that each $\mathcal{M}^{j}$ is an $\ell$-gapped family (this is trivial if $\ell=1$ since every family is 1-gapped, and it holds in Case 2 because $\mathcal{F}^{\prime}$ only contains elements who sizes are either even or odd).
We begin by describing the embedding process for $P$, ensuring that at every step we have the expected number of options for where to embed a given member. Let $H_{j}(P):=H(P)[\{v_{1},v_{2},v_{j}\}]$. For this, we will iteratively find partial embeddings $\varphi^{j}:\{v_{1},v_{2},\dots,v_{j}\}\rightarrow\mathcal{F}^{\prime}$ with the following properties.
  1. $(C_{0})$:
    We have $\varphi^{1}(v_{1})\in\mathcal{L}^{I(v_{1})}(\mathcal{M}^{|P|})$.
@@ 2. $(C_{1})$:
    For all edges $v_{a}v_{b}$ in $H_{j}(P)$, there exists a $(\chi,Q)$ in $\mathcal{M}^{|P|-j}$ with $\varphi^{j}(v_{a})$ in the $I(v_{a})$th position of $Q$ and $\varphi^{j}(v_{b})$ in the $I(v_{b})$th position of $Q$.
@@ 3. $(C_{2})$:
    For all noncomparable pairs $v_{a},v_{b}\in\{v_{1},\dots,v_{j}\}$, $\varphi^{j}(v_{a})$ and $\varphi^{j}(v_{b})$ are not comparable in $\mathcal{B}_{n}$.
Note in particular, $(C_{0})$ and $(C_{1})$ ensures $\varphi^{j}(v_{c})\in\mathcal{L}^{I(v_{c})}(\mathcal{M}^{|P|-j})$ for all $v_{c}\in\{v_{1},\dots v_{j}\}$.
Furthermore, note that $(C_{1}),(C_{2})$ ensures the poset induced by $\varphi^{j+1}(v_{1}),\varphi^{j+1}(v_{2}),\dots,\varphi^{j+1}(v_{j+1})$ has Hasse diagram $H_{j+1}(P)$. As two posets are isomorphic if and only if their Hasse diagrams are isomorphic, we have that this produces induced copies of $P$. As such, it will suffice to show that there are many ways of completing our embedding while satisfying these properties.
We now inductively construct our embeddings and take note of how many choices we have at each step of our process. In the base case, we embed $\varphi^{1}(v_{1})\in\mathcal{L}^{I(v_{1})}(\mathcal{M}^{|P|})$, noting that this satisfies all of the $(C_{i})$ conditions vacuously. Because we constructed $\mathcal{M}^{|P|}$ using Theorem 5.9, we have that $|\mathcal{L}^{I(x_{1})}(\mathcal{M}^{|P|})|=\Omega(\min_{F\in\mathcal{F}^{\prime}}{n\choose|F|})$. Thus by definition of our cases, we have in Case 1 that
$|\mathcal{L}^{I(x_{1})}(\mathcal{M}^{|P|})|=\Omega\left(\binom{n}{n/2}\right)$
where we used that $\binom{n}{|F|}\geq\frac{\varepsilon}{4q}\binom{n}{n/2}$, and in Case 2 we have
$|\mathcal{L}^{I(x_{1})}(\mathcal{M}^{|P|})|=\Omega\left(n^{-2}{n\choose n/2}\right)$
where we used that $\binom{n}{|F|}=\Omega\left(\frac{1}{n^{2}}\binom{n}{n/2}\right)$, using (1). In Case 2, we will make up this loss of a factor of $n^{2}$ by using the fact our Case 2 family is $2$-gapped.
Now let $j\geq 1$ and suppose $(C_{1})$ and $(C_{2})$ hold for $\varphi^{j}$. We want to extend $\varphi^{j}$ to a partial embedding $\varphi^{j+1}$ so that $(C_{1})$ and $(C_{2})$ both hold for $\varphi^{j+1}$.
For notational convenience let $x:=v_{j+1}$, and $y$ denote the unique parent of $x$ which has already been embedded. We wish to find $\gamma n^{\ell|(I(x)-I(y))|}$ many choices to embed $x$. For further convenience, we only consider the case $y\geq x$, the other case being analogous.
Let $\mathcal{P}_{j}=\{\varphi^{j}(z):z\not\geq y,z\in\{v_{1},v_{2},\dots,v_{j}\}\}$ The following set $\mathcal{A}$ is the set of choices for where we wish to embed $x$
$\mathcal{A}=\mathcal{L}^{I(x)}(\mathcal{M}^{|P|-j-1}(\varphi^{j}(y),I(y)))-D^{*}(\varphi^{j}(y),\mathcal{P}_{j}),$
i.e. this is the set of $F$ which are not in the forbidden neighborhood of $\mathcal{P}_{j}$ and which are in the $I(x)$th position of some $Q$-marked chain of $\mathcal{M}^{|P|-j-1}$ which has $\varphi^{j}(y)$ in the $I(y)$th position.
We will now show that the number of choices is large.
@@
$|\mathcal{A}|\geq\gamma n^{\ell(I(y)-I(x))}$

### v2

@@ 4.3 Tools for Induced Posets
Let $\mathcal{M}$ be a family of $q$-marked chains. Let $i\in[q]$, $F\in\mathcal{B}_{n}$. Recall that for each $\chi\in\mathcal{C}$, we let $\mathcal{M}(\chi,F,i)$ denote the set of all $(\chi,Q)\in\mathcal{M}$ such that $F$ is the $i$-th member of $Q$, and we let $\mathcal{M}(F,i)=\bigcup_{\chi}{\mathcal{M}(\chi,F,i)}$.
We say a family $\mathcal{F}\subseteq\mathcal{B}_{n}$ is $\ell$-gapped if for every $F,G\in\mathcal{F}$, with $F\subsetneq G$, $|G-F|\geq\ell$. For $q\geq 1$, we say that a $q$-marked chain family $\mathcal{M}$ with markers from $\mathcal{F}$ is $\ell$-gapped if $\mathcal{F}$ is $\ell$-gapped.
Definition 4.18.
@@
For every $\delta>0$, and every $\ell,q$ positive integers, there exists $\gamma=\gamma(\delta,\ell,q)$ such that the following holds for every $i,s\in[q]$ such that $i\neq s$. Let $\mathcal{M}$ be a $\ell$-gapped family of $q$-marked chains with markers from ~$\widetilde{\mathcal{B}}_{n}$ and let $F\in\mathcal{L}^{i}(\mathcal{M})$. Suppose $F$ is $(i,s,\gamma)$-bad with respect to $\mathcal{M}$ and that $n$ is sufficiently large.
@@
$\operatorname{Prob}[\chi\cap\mathcal{W}_{2}\neq\emptyset|F\in\chi]$ $\leq\sum_{D\in\mathcal{W}_{2}}\frac{1}{\binom{|F|}{|F-D|}}$ by (5)
@@
$\leq\frac{\delta}{2},$
by the choice of $\gamma=\frac{\delta}{2(3\ell q)^{\ell q}}$. On the other hand by Proposition 4.16,
@@
$\operatorname{Prob}[\chi\cap\mathcal{W}\neq\emptyset|F\in\chi]\leq\delta.$
@@
5 Embedding Tree Posets
In this section we use the tools developed in the previous section to prove our two main supersaturation results: Theorem 1.3 and a balanced supersaturation version of Corollary 1.4 stated formally as Theorem 5.6. Both of these proofs will follow essentially the same scheme for embedding a tree poset $P$ into a large family $\mathcal{F}$ one member at a time by using Theorem 4.15 to show that at every step, we always have many choices for how to embed the next member of $P$ even when we forbid some number of “bad” choices in $\mathcal{F}$.
5.1 Embedding Process
We establish the following general embedding result Theorem 5.2 using Theorem 4.15. It basically states that given some sufficiently sparse forbidden set $\Gamma\subseteq 2^{\mathcal{F}}$, we can grow many induced copies of $P$ not in $\Gamma$. For the non-balanced supersaturation we actually do no need to have such a forbidden set (i.e. $\Gamma=\emptyset$) but for the balanced supersaturation result, $\Gamma$ will be the upward closure of all the sets which are already contained in too many induced copies (i.e. saturated sets) of $P$ in the current collection.
Definition 5.1.
Given an $\ell$-gapped family $\mathcal{F}$, we say a collection $\Gamma\subseteq 2^{\mathcal{F}}$ is a $(\gamma,\ell,\mathcal{F})$-bounded family if it satisfies the two properties below
  1. 1.
    We have $\{F\}\notin\Gamma$ for all $F\in\mathcal{F}$.
  2. 2.
    If $\mathcal{D}\not\in\Gamma$, then the number of $F\in\mathcal{F}$ such that $\{F\}\cup\mathcal{D}\in\Gamma$ is less than $\gamma n^{\ell}$.
@@ 5.1 Embedding Process
Let $P$ be a tree poset of height $k$. Let $q\geq k$ and $\ell$ be positive integers and $\varepsilon>0$ then there exists a constant $\gamma=\gamma(\varepsilon,q,\ell)$ such that the following holds for any $v\in P$ and for any $\mathrm{r}:P\to[q]$ a poset homomorphism. Let ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ be an $\ell$-gapped family and has Lubell weight $\mu(F)\geq q-1+\varepsilon$ and $\Gamma$ a $(\frac{\gamma}{2},\ell,\mathcal{F})$- bounded family. Then, there is $F\in\mathcal{F}$ such that the number of injective induced homomorphisms $\varphi:P\rightarrow\mathcal{F}^{\prime}$ satisfying $\varphi(P)\not\in\Gamma$ and $\varphi(v)=F$ is at least
$\left(\frac{\gamma}{2}\right)^{|P|-1}\left(\prod_{xy\in H(P)}n^{\ell|\mathrm{r}(y)-\mathrm{r}(x)|}\right).$
@@
Because $P$ is a tree poset, there exists an ordering $v_{1},v_{2},\dots v_{|P|}$ of the members of $P$ with $v_{1}=v$ such that every member $v_{j}$ with $j\geq 2$ has exactly one neighbor $v_{j^{\prime}}$ with $j^{\prime}<j$ in the Hasse diagram of $P$, and we call this $v_{j^{\prime}}$ the parent of $v_{j}$.
Apply Theorem 4.15 to $\mathcal{F}$ with $q,\varepsilon$ to find a $\delta=\delta(\varepsilon,q)>0$ and a nested sequence of $q$-marked chains $\mathcal{M}^{|P|}\subseteq\mathcal{M}^{|P|-1}\subseteq\cdots\subseteq\mathcal{M}^{0}$ with markers from $\mathcal{F}$ such that for all $j\in[|P|]$ we have that for each $(\chi,Q)\in\mathcal{M}^{j}$ and $F\in Q$, we have that $F$ is $\delta$-robust with respect to $\mathcal{M}^{j-1}$.
Fix $\gamma$ be obtained from Lemma 4.20 applied with $q,\ell,\delta$.
Fix $F_{0}\in\mathcal{L}^{\mathrm{r}(v_{1})}(\mathcal{M}^{|P|})$ and define our initial embedding $\varphi^{1}(v_{1}):=F_{0}$. To define the final embedding, we iteratively extend $\varphi^{j}:\{v_{1},v_{2},\dots,v_{j}\}\rightarrow\mathcal{F}$ to $\varphi^{j+1}$ maintaining the following properties.
@@ 1. $(C_{1})$:
    Let $H_{j}(P):=H(P)[\{v_{1},v_{2},\ldots,v_{j}\}]$, that is the Hasse diagram induced by the first $j$ elements of $P$. For all edges $v_{a}v_{b}$ in $H_{j}(P)$, there exists a $(\chi,Q)$ in $\mathcal{M}^{|P|-j}$ with $\varphi^{j}(v_{a})$ in the $\mathrm{r}(v_{a})$th position of $Q$ and $\varphi^{j}(v_{b})$ in the $\mathrm{r}(v_{b})$th position of $Q$.
@@ 2. $(C_{2})$:
    For all noncomparable pairs $v_{a},v_{b}\in\{v_{1},\dots,v_{j}\}^{2}$, $\varphi^{j}(v_{a})$ and $\varphi^{j}(v_{b})$ are not comparable in $\mathcal{B}_{n}$.
  3. $(C_{3})$:
    $\{\varphi^{j}(v_{1}),\varphi^{j}(v_{2}),\dots\varphi^{j}(v_{j})\}\not\in\Gamma$.
Note that these properties are satisfied for $\varphi^{1}$, since $(C_{1}),(C_{2})$ both hold vacuously and $(C_{3})$ holds because $F_{0}\not\in\Gamma$ by Definition 5.1 part 1.
Let us emphasize that the choice of $F_{0}$ and $(C_{1})$ ensures $\varphi^{j}(v_{c})\in\mathcal{L}^{\mathrm{r}(v_{c})}(\mathcal{M}^{|P|-j})$ for all $v_{c}\in\{v_{1},\dots v_{j}\}$. This will be useful later.
Furthermore, $(C_{1}),(C_{2})$ ensures the poset induced by $\varphi^{j+1}(v_{1}),\varphi^{j}(v_{2}),\dots,\varphi^{j}(v_{j})$ has Hasse diagram isomorphic to $H_{j}(P)$. As two posets are isomorphic if and only if their Hasse diagrams are isomorphic, by the end of the process we will obtain an induced copy $P^{\prime}$ of $P$. Let us now show that indeed these partial embeddings are possible to construct, and count how many choices we have at each step.
Now let $j\geq 1$ and suppose $(C_{1}),(C_{2}),$ and $(C_{3})$ hold for $\varphi^{j}$. We want to extend $\varphi^{j}$ to a partial embedding $\varphi^{j+1}$ so that $(C_{1})-(C_{3})$ hold for $\varphi^{j+1}$.
For notational convenience let $x:=v_{j+1}$, and let $y$ denote the unique parent of $x$ which has already been embedded. We wish to find at least one way of embedding $x$. For further convenience, we only consider the case $y>_{P}x$, the other case being analogous. This in particular implies $\mathrm{r}(y)<\mathrm{r}(x)$.
Recall the formal definition of $D^{*}(F,\mathcal{S})$ from Section 4.3 for a member $F\in\mathcal{F}$ and a family $\mathcal{S}$ which is simply the set of all members $S\in\mathcal{S}$ which are downsets of $F$ and are comparable to some set of $\mathcal{S}$. Furthermore, let ^$\widehat{\mathcal{M}}^{|P|-j-1}:=\mathcal{M}^{|P|-j-1}(\varphi^{j}(y),\mathrm{r}(y))$, that is the family of all $q$-marked chains in $\mathcal{M}^{|P|-j-1}$ which have $\varphi^{j}(y)$ in their $\mathrm{r}(y)$th position. Let $\mathcal{P}_{j}=\{\varphi^{j}(z):z\not\geq_{P}y,z\in\{v_{1},v_{2},\dots,v_{j}\}\}$, that is the set of all images of currently embedded elements of $P$ which are either below $y$ or incomparable with $y$, and thus not cannot be comparable with $\varphi^{j+1}(x)$ because in the embedding $\varphi^{j}$, $x$ has the unique neighbor $y$.
The following set encodes the choices for $x$ that would preserve $\varphi^{j+1}$ being an induced homomorphism. Now let
$\mathcal{A}=\mathcal{L}^{\mathrm{r}(x)}(\widehat{\mathcal{M}}^{|P|-j-1})-D^{*}(\varphi^{j}(y),\mathcal{P}_{j}),$
i.e. this is the set of $F$ which are not in the forbidden neighborhood of $\mathcal{P}_{j}$ and which are in the $\mathrm{r}(x)$th position of some $q$-marked chain of $\mathcal{M}^{|P|-j-1}$ which has $\varphi^{j}(y)$ in the $\mathrm{r}(y)$th position.
@@ 3. $(C_{3})$:
$|\mathcal{A}|\geq\gamma n^{\ell(\mathrm{r}(x)-\mathrm{r}(y))}$
Proof of Claim 5.3.
Suppose $|\mathcal{A}|<\gamma n^{\ell(\mathrm{r}(x)-\mathrm{r}(y))}$. We derive a contradiction by showing this would imply $\varphi^{j}(y)$ is $(\mathrm{r}(y),\mathrm{r}(x),\gamma)$-bad (recall Definition 4.18) with $\mathcal{W}_{1}=\mathcal{P}_{j}$ and $\mathcal{W}_{2}=\mathcal{A}$. This is a contradiction because Lemma 4.20 implies that $\varphi^{j}(y)$ is $(\mathrm{r}(y),\delta)$-bad with respect to $\mathcal{M}^{|P|-j-1}$. However, this cannot happen since we constructed our marked chain families using Theorem 4.15 which guarantees that $\varphi^{j}(y)$ is $\delta$-robust with respect to $\mathcal{M}^{|P|-j-1}$ (see Definition 4.6).
Note that by definition of $\mathcal{P}_{j}$ we have $\mathcal{P}_{j}\cap U(\varphi^{j}(y))=\emptyset$ and $|\mathcal{P}_{j}|\leq|P|$, and by hypothesis we are assuming $|\mathcal{W}_{2}|\leq\gamma n^{\ell(\mathrm{r}(x)-\mathrm{r}(y))}$. It thus remains to check the last condition in the definition of $(\mathrm{r}(y),\mathrm{r}(x),\gamma)$-badness, i.e. that for each $(\chi,Q)\in\widehat{M}^{|P|-j-1}$, either $Q\cap D^{*}(\varphi^{j}(y),\mathcal{P}_{j})\neq\emptyset$ or $Q$ contains a member from $\mathcal{A}$ in the $\mathrm{r}(x)$th position.
Consider any $(\chi,Q)\in\widehat{\mathcal{M}}^{|P|-j-1}$. If $Q\cap D^{*}(\varphi^{j}(y),\mathcal{P}_{j})=\emptyset$, then the $\mathrm{r}(x)$th member of $Q$ lies in $\mathcal{A}$ by definition of $\mathcal{A}$. Therefore, every $(\chi,Q)\in\widehat{\mathcal{M}}^{|P|-j-1}$ has $Q$ either intersecting $D^{*}(\varphi^{j}(y),\mathcal{P}_{j})$ or $\mathcal{A}$. ∎
Let $\mathcal{A}^{\prime}=\mathcal{A}-\{F\in\mathcal{F}:\{F,\varphi^{j}(v_{1}),\varphi^{j}(v_{2}),\dots\varphi^{j}(v_{j})\}\in\Gamma\}$. By $(C_{3})$ and Definition 5.1, $|\mathcal{A}^{\prime}|\geq\frac{1}{2}\gamma n^{\ell(\mathrm{r}(x)-\mathrm{r}(y))}$. Fix one such $F\in\mathcal{A}^{\prime}$. We claim that by embedding $x$ into any set in $\mathcal{A}^{\prime}$, we will extend $\varphi^{j}$ to $\varphi^{j+1}$, satisfying all the desired properties $(C_{1})-(C_{3})$. Indeed, define $\varphi^{j+1}(v_{a})=\varphi^{j}(v_{a})$ for all $a\in[j]$ and $\varphi^{j+1}(x)=F$. Let us now check that $\varphi^{j+1}$ satisfies $(C_{1})-(C_{3})$.
Since $\mathcal{M}^{|P|-j}\subseteq\mathcal{M}^{|P|-j-1}$, for all $a,b\in[j]$, $(C_{1})$ and $(C_{2})$ hold for all such pairs so we only need to check that these two conditions are satisfied for pairs of form $(x,v_{a})$ for any $a\in[j]$.
By definition of $\mathcal{A}$, $\varphi^{j+1}(x)$ is in the $\mathrm{r}(x)$th position of some $(\chi,Q)\in\mathcal{M}^{|P|-j-1}$ which has $\varphi^{j+1}(y)$ in the $\mathrm{r}(y)$th position.
Let $a\in[j]$, and suppose $v_{a}$ is not comparable with $x$. We would like to show that $\varphi^{j+1}(v_{a})$ and $F$ are incomparable. It is easy to see that $v_{a}\not\geq y$, so $\varphi^{j+1}(v_{a})=\varphi^{j}(v_{a})\in\mathcal{P}_{j}$. Since $F\in D(\varphi^{j}(y))\setminus D^{*}(\varphi^{j}(y),\mathcal{P}_{j})$ by our choice, $F$ is not comparable with any element in $\mathcal{P}_{j}$, in particular, $\varphi^{j+1}(v_{a})$.
To see $\varphi^{j+1}$ satisfies $(C_{3})$, notice that $F$ was chosen so that $\{F,\varphi^{j}(v_{1}),\varphi^{j}(v_{2}),\dots\varphi^{j}(v_{j})\}\not\in\Gamma$.
Finally, to check that the desired counting bound does hold, it is enough to observe that at every step while going from $\varphi^{j}$ to $\varphi^{j+1}$ starting at $j\geq 2$ we had at least $\frac{1}{2}\gamma n^{\ell(\mathrm{r}(x)-\mathrm{r}(y))}$ many choices to embed $x=v_{j+1}$, where $y$ is the parent of $x$ among already embedded vertices $v_{1},v_{2},\dots,v_{j}$. Taking the product over all $j=2,\dots,|P|-1$ gives us the desired result. ∎
5.2 Proof of Theorem 1.3
We begin by proving an analogue of Theorem 1.3 in the setting of Lubell weight and $\ell$-gapped families and deduce Theorem 1.3 from it.
Theorem 5.4.
For every tree poset $P$ of height $k$, positive integers $q,\ell$ with $q\geq k$, and $\varepsilon>0$, there exists $\gamma=\gamma(\varepsilon,q,\ell)$ such that the following holds for any $\mathrm{r}:P\rightarrow[q]$ a poset homomorphism. Let ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ be an $\ell$-gapped family such that for all $\mathcal{S}\subseteq\mathcal{F}$ with $|\mathcal{S}|=N$ we have $\mu(\mathcal{F}-\mathcal{S})\geq q-1+\varepsilon$. Then, the number of induced copies of $P$ in $\mathcal{F}$ is at least
$\left(\frac{\gamma}{2}\right)^{|P|-1}\left(\prod_{xy\in H(P)}n^{\ell|\mathrm{r}(y)-\mathrm{r}(x)|}\right)\cdot N.$